OpenStudy (anonymous):

what is the sum of the infinite arithmetic series? (-5)+0+5+10+...+65 a.900 b.455 c.450 d.445

OpenStudy (anonymous):

This is not an infinite series...

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

No, it's not. It's finite. If \(a_n\) gives the \(n\)th term in the series, then \[a_1=-5\\ a_2=-5+5=0\\ a_3=-5+5(2)=5\\ \quad~~\vdots\\ a_n=-5+5(n-1)\] From the last equation, you can determine how many terms are in the series. \[65=-5+5(n-1)\] Clearly \(n\) is a finite number.